OpenStudy (anonymous):

Karen wants to distract her noisy four-year-old child at a restaurant. She pulls out a bag that has 4 black marbles, 8 red marbles, and 9 green marbles. The child likes green marbles. What is the probability Karen will randomly pick a green marble from the bag in one try?

Mehek (mehek14):

green marble = 9 total # of marbles = 21 set up a fraction \(\dfrac{9}{21}\) divide by 3 \(\dfrac{3}{7}\)

OpenStudy (wil476003):

43 percent or 9/21 3/7

OpenStudy (anonymous):

Total no.of marbles = (4+8+9) = 21 No.of green marbles = 9 therefore, P(green) = 9/21 dividing above and below by 3 we get: 3/7

OpenStudy (anonymous):

thx